OpenStudy (anonymous):

Write 3.1818…. as a fraction in simplest form.

OpenStudy (anonymous):

Write the decimal number as a fraction (over 1) 3.1818 = 3.1818 / 1 Multiplying by 1 to eliminate 4 decimal places we multiply top and bottom by 4 10's Numerator (N) N = 3.1818 × 10 × 10 × 10 × 10 = 31818 Denominator (D) D = 1 × 10 × 10 × 10 × 10 = 10000 N / D = 31818 / 10000 Simplifying your fraction = 31818/10000 = 15909/5000 = 3 909/5000 hope i helped

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

lol she made that look complicated O.o but gud explanation :)

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

next question

OpenStudy (anonymous):

3. Use a factor tree to write the prime factorization of 430.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The Prime Factors are: 2 x 5 x 43 In Exponential Form: 21 x 51 x 431 430 / / 2 215 / / 5 43

OpenStudy (anonymous):

wait i forgot to put the ^ in the exponents

OpenStudy (anonymous):

here it is again 2^1 * 5^1 * 43^1

OpenStudy (anonymous):

wow you r good DD

OpenStudy (anonymous):

thanks hope i helped u

OpenStudy (anonymous):

next question anyone

OpenStudy (anonymous):

4. Find the LCM of 15 and 27.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

135 is the lcm you list the multiples of both numbers the 1st common number you see in both list of number is the least common mutliply

OpenStudy (anonymous):

are you a teacher or sum cuz you are smart

OpenStudy (anonymous):

no im a 7th grade student im in honors class

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

wat grade r u in\

OpenStudy (anonymous):

5. Subtract 6 ¾ - 4¹¹⁄₁₂. Write the answer in simplest form.

OpenStudy (anonymous):

okay but may i ask wat grade ur in

OpenStudy (anonymous):

12 but im bored :P

OpenStudy (anonymous):

ooh wow ur like in college

OpenStudy (anonymous):

wat college do u go to

OpenStudy (anonymous):

yepp

OpenStudy (anonymous):

1 5/6

OpenStudy (anonymous):

6. Find the perimeter

OpenStudy (anonymous):

can i answer it later im at skool n i have to put away the computer

OpenStudy (anonymous):



OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i bet you mean the repeating decimal

OpenStudy (anonymous):

\[3.1818…\] with the \(.18\) repeating

OpenStudy (anonymous):

nope Find the perimeter

OpenStudy (anonymous):

Write 3.1818…. as a fraction in simplest form

OpenStudy (anonymous):

it means the \(.18\) repeats

OpenStudy (anonymous):

\[.181818...=\overline{.18}=\frac{18}{99}=\frac{2}{11}\]

OpenStudy (anonymous):

and while @arianadarsh gave a good explanation, it was an answer to a different question

OpenStudy (anonymous):

2. Find three consecutive even integers whose sum is 180